Use a specified method to solve the system of equations.
(graphically)
y=4-x
y=x-6
I understand that the first ordered pair is (-1,4) and the second is (1,-6)
I just have no clue how to graph these.

Answer by mananth(16946)

y=4-x
y=x-6
Find the x & y intercept points by plugging x=0,y=0 in both equations.
equation I
when x= 0 y= 4 ,( 0 , 4 )
when y = 0 x= 4 ,( 4 , 0 )
Equation 2
when x= 0 y= -6 ( 0 , -6 )
when y = 0 x= 6 ( 6 , 0 )



You will observe that the point of intersection is 5,-1)
The solution of the system is ( 5 ,-1 )